A lady uses + 1.5 D glasses to have normal vision from 25 cm onwards. She uses a 20 D lens as a simple microscope to see an object. Find the maximum magnifying power if she uses the microscope (a) together with her glass (b) without the glass. Do the answers suggest that an object can be more clearly seen through a microscope without using the correcting glasses?

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. The lady uses +1.5D glasses to have normal vision at 25 cm. So, with the glasses, her least distance of clear vision = D = 25 cm Focal length of the glasses = 1/1.5 m = 100/1.5 cm So, without the glasses her least distance of distinct vision should be more If, u = - 25cm, f = 100/1.5 cm Now, 1/v – 1/u = 1/f = 1.5/100 – 1/25 = 1.5 – 4/100 = -235/100 ⇒ v = - 40cm = near point without glasses. Focal length of magnifying glass = 1/20 m = 0.05m = 5 cm = f (a) The maximum magnifying power with glasses m = 1 + D/f = 1 + 25/5 = 6 [∵ D = 25cm] (b) without the glasses, D = 40cm So, m = 1 + D/f = 1 + 40/5 = 9